OpenStudy (anonymous):

Express 4xi + 5yi^8 + 6xi^3 + 2yi^4 in simplest a+bi form

OpenStudy (anonymous):

\[4x \iota+5y \iota^8+6x \iota^3+2y \iota^4\] \[=4x \iota+5y \left( \iota^2 \right)^4+6x \left( \iota^2 \right)\iota+2 y \left( \iota^2 \right)^2\] \[\iota^2=-1\] simplify it.

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 4xi\implies 4x\cdot i \\ \quad \\ 5yi^8\implies 5y\cdot i^2\cdot i^2\cdot i^2\cdot i^2 \\ \quad \\ 6xi^3\implies 6x\cdot i^2\cdot i \\ \quad \\ 2yi^4 \implies 2y\cdot i^2\cdot i^2 \end{array} \\ \quad \\ \bf {\color{blue}{ recall\implies i^2\implies \sqrt{-1}\cdot \sqrt{-1}\implies \sqrt{(-1)^2}\implies -1}}\) what do you think?